You are the media planning manager in a large online firm’s marketing department. The budget for your new product’s upcoming TV ad campaign is $3.5 million and the cost per thousand exposures in the market you are targeting is $12.50. The average exposure frequency is 14. Out of the viewers you reach, 2.8% will go online to search for your product on your website, and out of these 19% will actually buy your product at an average price of $64.75. Your target for this campaign is an online sales volume of $7 million. Is your advertising budget sufficient to make this happen? If so, how much do you have left over? If not, how much more would you need to spend?
Answer #1

To achieve a sale of $7 million or $700000

the  number of sales required @ $64.75 will be

700000/64.75= 10811

since only 19% of on line search customers actually purchase the product so the required number of on line customer for 10811 sale will be 10811*100/19= 56900

Now only 2.8% of viewers of ad will be on line for search so the required number of viewers of ad for 56900 on line customer will be 56900*100/2.8= 2032143

average exposure frequency is 14 so we need 232143*14= 28450002 exposures

For 1000 exposures we $12.5 so for 28450002 exposures we need 28450002*12.5/1000 $355625

and the fund is $3.5 million or $350000

therefore we are short by $355625-$350000= $5625